# references.bib

### v2

@@ Methods
    During $[0,t_{1})$, the representation is $\vec{V}_{F}=V_{1}$ and $\vec{V}_{J}=V_{2}$. Reactions: $\mathcal{S}=\left\{\mathcal{R}_{2},\mathcal{R}_{3}\right\}$.
@@
    During $[t_{1},t_{2})$, the representation is $\vec{V}_{F}=(V_{1},V_{2})^{\intercal}$ and $\vec{V}_{J}=\varnothing$. Reactions: $\mathcal{S}=\varnothing$ (ODE regime).
@@
    During $[t_{2},t_{3})$, the representation is $\vec{V}_{F}=V_{2}$ and $\vec{V}_{J}=V_{1}$. Reactions: $\mathcal{S}=\left\{\mathcal{R}_{1},\mathcal{R}_{3}\right\}$.
@@
    From $t_{3}$ onwards the representation is $\vec{V}_{F}=\varnothing$ and $\vec{V}_{J}=(V_{1},V_{2})^{\intercal}$. Reactions: $\mathcal{S}=\mathcal{R}$ (CTMC regime).
@@
Figs. 1D and E show random trajectories sampled from the Lotka-Volterra process as represented with JSF. Fig. 1D shows several cycles of the two populations and Fig. 1E demonstrates how, with the inclusion of discrete stochastic behaviour, the predator species can go extinct, allowing the prey to grow exponentially. Full details, including description of the algorithms used to sample from the JSF process are given in SI Sections 1 and 2.
Data and Software Availability
All data and code to reproduce our results is available at https://github.com/DGermano8/JSFGermano2024. A Python package implementing the Jump-Switch-Flow method is available at https://dgermano8.github.io/JSF.
Figure 1: JSF provides a way to capture both the continuous deterministic dynamics of large populations and the discrete stochastic dynamics of small populations. A A compartmental model representation of the Lotka-Volterra system ((2)): prey (which reproduce at rate $\alpha$), predators (which die at rate $\gamma$), and their interaction (predating at a rate proportional to $\beta$). B The compartmental model can be formalised as a reaction network via a stoichiometry matrix $\eta$ which details each species role in the reactions. The reactants consumed $\eta^{-}$ and the products produced $\eta^{+}$ can be written explicitly with $\eta=\eta^{+}-\eta^{-}$. C A cartoon example demonstrating how the representation captures the continuous variation of some compartment-reaction pairs (flow) and the discrete stochastic changes (jumps) in others. Transitioning (switching) between a continuous and discrete state occurs at a threshold $\Omega$ which is indicated with a horizontal dashed line. D An example trajectory from the Lotka-Volterra model exhibiting the characteristic cyclical behaviour with additional domesticity influencing the dynamics at low population sizes. E An example trajectory showing the possibility for the predator species to go extinct and subsequent growth of the prey species.
@@ Results
The SIRS model with demography is an extension of the classic susceptible-infectious-removed (SIR) model. In the SIR model susceptible individuals may be infected by contact with infectious individuals; infected individuals eventually cease to be infectious and transition to the removed compartment. The SIRS model with demography extends the SIR model by allowing removed individuals to transition back to being susceptible to infection and for individuals to give birth to new (susceptible individuals) and die (at equal rates). Fig. 2A shows a compartmental diagram of the SIRS model with demography.
Fig. 2B shows our simulated (via the Doob-Gillespie algorithm) time series of daily noisy measurements of the prevalence of infection. The parameters used in the simulation are shown in Fig. 2C; these values are broadly consistent with existing estimates for influenza or SARS-CoV-2. We assume that these measurements are drawn from a negative binomial distribution where the expected value is equal to the true prevalence of infection and there is a constant (known) dispersion parameter, $k=100$.
@@
Combining the SIRS model with a particle filter (Moss2024; kitagawa1996monte) we used the first 100 days of the time series to forecast the remaining 150 days. Fig. 2B shows the estimated true prevalence across the first 100 days and forecasts the prevalence for the subsequent 150 days. Included in the forecast is a daily estimate of the probability that the pathogen has been eliminated (in the simulation the pathogen was eliminated on day 250). It is important to note that there is an endemic equilibrium in this model, so it is reasonable for the probability to plateau as it does.
To demonstrate the utility of the posterior distribution (i.e. the capacity of the particle filter to forecast the epidemic after the 100 days of observed data), we considered the impact of a possible intervention. This intervention, introduced on day 100, reduces the force of infection (and hence the effective reproduction number) by a factor of $\alpha$. Fig. 2D shows how the probability of eliminating the pathogen increases substantially as we decrease $\alpha$ from $1$ (no intervention) to $0.7$ (a $30\%$ reduction in transmission).
@@
Figure 2: JSF enables us to forecast when a pathogen will be eliminated from a population, i.e. when we can claim that an epidemic has “ended”, and quantify the likely impact of varying levels of intervention. A A compartmental model representation of the SIRS with demography. B A simulated time series of noisy measurements of prevalence along with inferred prevalence trajectories and forecasts of future prevalence with associated uncertainty bands. The dashed vertical line indicates the date up until which we have data. The solid blue line indicates an estimate of the probability that the pathogen has been eliminated from the population (as opposed to persisting at very low levels). C The true parameters used in the simulation along with their posterior estimates and the associated priors used. Each of the marginal posterior distributions peaks tightly about the true value. D The probability of pathogen elimination increases with the strength of the intervention against transmission. Near certainty of elimination is achieved with a $30\%$ reduction of transmission.
@@
To further investigate the computational properties of JSF and to compare it with exact (Doob-Gillespie) simulation and (approximate) tau-leaping simulation we used repeated simulations from the SIRS model with demography. We partitioned the simulations into three classes: early stochastic extinction, fade-out after a single epidemic, and sustained transmission (SI Section 3.) As seen in Table S3, JSF and exact simulation generate very similar proportions of samples in each of these classes. Two aspects of this system of practical importance are the timing and magnitude of peak infections, and the total number of infections. As can be seen in Fig. S4 the distribution of peak timing and total number of infections are in good agreement between JSF and exact simulation. For the magnitude of peak infections, there is substantially less variability in the JSF samples than the exact samples, but the proportional difference is small.
With regards to computational efficiency, Fig. S7 shows the average time required for simulations using each of JSF, exact simulation and the tau-leaping approximation. We see that for populations of size above about $10^{5}$, (i.e. approximately the size of a small city) JSF is substantially faster than both exact simulation and the tau-leaping approximation.
@@
Figure 3: The TEIRV model describes the within-host dynamics of host cells and virus during infection. The model accounts for delays in virus production post cellular infection and the capability of target cells to enter a refractory state as a defence against infection. A A compartmental diagram of the TEIRV model showing the interactions between: Target cells ($T$), Refractory cells ($R$), cells in the Eclipsed phase of infection ($E$) and Infectious cells ($I$); and the Virions ($V$). Solid arrows represent the exchange of mass (cells or virions) through compartments, while dashed lines represent the influence of a compartment on a reaction’s rate of occurrence. B The representation of this model as a reaction network with a stoichiometric matrix. This formalises the dependency on different variables indicated by the dashed arrows in the compartmental diagram. C An example trajectory showing the exponential growth and decline in the amount of free virus across the duration of the infection, this process terminates in the virus, eclispse and infected cell populations reaching zero (shortly before time $12$). Reaching zero which is possible in this hybrid continuous/discrete stochastic model but does not occur in a purely ODE based representation.
Understanding viral clearance is important when deciding how long treatment must be maintained. This case study demonstrates how JSF enables inference of virus clearance, which hitherto has been computationally challenging yan2016extinction; farrukee2018characterization. In particular, we use a mathematical model — the TEIRV model described below — to study viral clearance using longitudinal data from 60 individuals infected with SARS-CoV-2 ke2022daily.
@@
The TEIRV model is an extension of the classic target-infected-virus (TIV) model perelson2002modelling. In the TIV model target cells may be infected by the virus; before dying, infected cells produce virus; and the virus can degrade or infect remaining target cells. The TEIRV extends the TIV model through the inclusion of an eclipsed compartment, to model the delay between infection of a target cell and the subsequent production of virus, and a refractory compartment, to model heightened antiviral defences of target cells, e.g. through the effects of interferons ke2022daily; blancomelo2020imbalanced. To simplify the use of this model a quasi-steady-state approximation for interferon production is employed. This approximation allows us to avoid the need to explicitly model the amount of interferon present.
Fig. 3A shows a compartmental diagram of the TEIRV model. Target cells becoming infected by virions at rate $\beta$, which then enter the eclipsed phase. These eclipsed cells then become infectious at rate $k$. Infectious cells are cleared from the population at rate $\delta$, and produce virions at rate $\pi$. Virions are themselves cleared at rate $c$. The infectious cells recruit interferon, which cause the target cells to become refractory (and hence protected against infection) at rate $\Phi$. The refractory cells return to a naive state as target cells at rate $\rho$. These assumptions are represented with the following ODE system:
@@
The stoichiometric matrix corresponding to these assumptions is shown in Fig. 3B. Fig. 3C, which shows a trajectory sampled from the TEIRV model when represented with JSF; the classic boom-bust dynamics of the viral populations can be seen in the exponential growth and decline of the $V$ compartment. Unlike solutions of the ODE model in (3), we see that by time $t=14$, the populations of eclipsed and infected cells, and the virus have gone extinct, indicating a definitive end to the infection.
The basic reproduction number, $\mathcal{R}_{0}$ is a fundamental quantity of epidemiological models. For the TEIRV model, $\mathcal{R}_{0}$ can be calculated from the next-generation matrix diekmann2010construction:
@@
The longitudinal data contains cycle number (CN) values from nasal samples. The CN values are inversely proportional to the number of virions present. In our analysis we used an existing (empirical) model to link the cycle numbers to the (logarithmic) viral genome load ke2022daily: $\log_{10}V=11.35-0.25\text{CN}$. We assume the observed values are drawn from a normal distribution with mean $\log_{10}V$ and unit standard deviation, and that the values are truncated at the detection limit of $-0.65$.

### v3

@@ Methods
    Before $t_{1}$, the representation is $\vec{V}_{F}=v_{1}$ and $\vec{V}_{J}=v_{2}$ and the reactions are $\mathcal{R}_{J}=\left\{{r}_{2},{r}_{3}\right\}$ (hybrid regime).
@@
    During $[t_{1},t_{2})$, the representation is $\vec{V}_{F}=(v_{1},v_{2})^{\intercal}$ and $\vec{V}_{J}=\varnothing$ and the reactions are $\mathcal{R}_{J}=\varnothing$ (ODE regime).
@@
    During $[t_{2},t_{3})$, the representation is $\vec{V}_{F}=v_{2}$ and $\vec{V}_{J}=v_{1}$ and the reactions are $\mathcal{R}_{J}=\left\{{r}_{1},{r}_{3}\right\}$ (hybrid regime).
@@
    From $t_{3}$ onwards the representation is $\vec{V}_{F}=\varnothing$ and $\vec{V}_{J}=(v_{1},v_{2})^{\intercal}$ and the reactions are $\mathcal{R}_{J}=\mathcal{R}$ (CTMC regime).
@@
Figures 1D and E show two instances of random trajectories sampled from the Lotka-Volterra process as represented with JSF with a switching threshold of $\Omega=30$. In the trajectory in Figure 1D the populations undergo several cycles while in Figure 1E, with the inclusion of discrete stochastic behaviour, the predator species goes extinct, allowing the prey to grow exponentially.
Mathematical detail
Here we provide the mathematical details of JSF: the evolution and computation of values for flowing compartments (see ‘Flow events’); the evolution and computation of values for jumping compartments (see ‘Jump events’); the sampling of jump times (see ‘Jump clock updates’); and the switching of compartments between jumping and flowing regimes (see ‘Switching events’). Further details, including pseudo-code listings of the algorithms used to sample from the JSF process, are given in SI Section 1.
Flow events
Equation (5) contains the differential equations describing the evolution of the flowing compartments. If we consider an interval of time between jumps, $\left[t_{0},t_{1}\right]$, then $\vec{V}_{J}(t)$ remains constant and therefore, $\vec{\Lambda}_{J}(t,\vec{V})=\vec{0}$ across this interval. Therefore, the differential equations for the flowing compartments are:
$\frac{\mathrm{d}\vec{V}_{F}}{\mathrm{d}t}=\eta_{FF}\vec{\lambda}_{F}(t,\vec{V}).$ (8)
This is a standard dynamical system of ODEs, which we numerically integrate forward in time over discrete time steps of size $\Delta t$. While an arbitrary order method may be used, we use a simple Forward Euler method:
$\vec{V}_{F}(t+\Delta t)=$ $\vec{V}_{F}(t)+\Delta\vec{V}_{F}(t)$ (9)
$=$ $\vec{V}_{F}(t)+\Delta t\,\eta_{FF}\vec{\lambda}_{F}(t,\vec{V})+O(\Delta t^{2}).$ (10)
Figures 2A and 2B depict examples of a flowing compartment and a jumping compartment, respectively. Flow updates are given at intervals of discrete time step $\Delta t$ as indicated. In the Jump-Switch-Flow algorithm, flowing time steps are bounded above by a finite, but small, $\Delta t$. Flowing time steps have this maximum duration if no jumps are found on this time interval (that is, if (8) remains valid). If a jump occurs at $\Delta\tau<\Delta t$ from any given time (see Section ‘Jump events’) then a flowing time step of $\Delta\tau$ is used instead of $\Delta t$ in Equation (9). In this case, a jump is instantaneously applied after the flow event and the next interval of time is associated with new flow rates $\vec{\lambda}_{F}$ (which are therefore updated at the exact moment of time of the jump event). We note therefore that error associated with the ODE solver is determined by the size of the finite capped time interval size $\Delta t$ according to classical ODE error analysis. In this case since we use Forward Euler, the error scale like $O(\Delta t^{2})$. Using Equation (9), over each time step (flowing interval), flow events only directly affect flowing compartments as can be seen in Figure 2A. Jumping compartments like that shown in Figure 2B are unaffected directly by flowing events because the moments of a jumping event are pre-determined according to approximations in the trajectories of the flowing compartments which have the same accuracy as the ODE solver (for example, we assume a linear changes in the flowing compartments as consistent with Forward Euler method used in Equation (9) – see Section ‘Jump events’ for details. As indicated by discontinuities in both Figure 2A and Figure 2B, jump events may be applied to flowing and jumping compartments alike in order to conserve mass in a manner which we shall now describe.
Figure 2: The JSF process can be numerically integrated. A and B: Dynamic changes in flowing compartments (A) and jumping compartments (B). Flows are computed using the time step $\Delta t$. If a jump occurs within a flow update, the flows are adjusted up until the corresponding time $\Delta\tau$ as a time step instead of $\Delta t$ (indicated by a discontinuity in both A and B). Jumps occur when a jump clock reaches zero. C: The jump clock $J_{k}$ for reaction $k$ counts down until event $k$ occurs when $J_{k}=0$, i.e. between $t$ when $J_{k}>0$ and $t+\Delta t$ when $J_{k}<0$. The time of the jump is found by interpolating for $t+\Delta\tau$ using Equation (18), where $\Delta\tau<\Delta t$. Once the jump has occurred and jumps implemented according to the respective stoichiometric updates, $J_{k}$ is reinitialised by sampling $J_{k}=-\log(u_{k})$ where $u_{k}\sim\text{Unif}(0,1)$, and then counts down with successive time steps $\Delta t$ until the next jump $k$ is realised. D A switching event of type 1 occurs when a discrete compartment jumps above the switching threshold, $\Omega$. This is shown here after the jump time step $\Delta\tau_{2}$. E A switching event of type 2 occurs when a flowing compartment reaches the switching threshold, $\Omega$. Here, the time at which the switch occurs is found using continuation. F A switching event of type 3 occurs when a flowing compartment is caused to go below the switching threshold, $\Omega$, due to a jump in another compartment. Here the jump has caused the flowing compartment to end at the dotted circle (which we denote as $\hat{v}_{i}$). This compartment is then re-labelled as jumping. Since jumping compartments require integer states, the state is re-initialised to take the value $\lceil\hat{v}_{i}\rceil$ with probability $v_{i}-\lfloor\hat{v}_{i}\rfloor$, and otherwise $\lfloor\hat{v}_{i}\rfloor$ such that the expected value of the state is at the dotted circle (in the example presented here, the variable was rounded down). For all switching events, membership of $V_{F}$ and $V_{J}$ is recomputed after switching.
Jump events
The reactions in $\mathcal{R}_{J}$ produce discontinuous jumps in the state vector $\vec{V}$. In Equations (5) and (6), each element in the vector $\vec{\Lambda}_{J}$ corresponds to a reaction in $\mathcal{R}_{J}$. Consider, for example, $r_{k}\in\mathcal{R}_{J}$. The $k^{\text{th}}$ element of $\vec{\Lambda}_{J}$ is:
$\Lambda_{J}^{(k)}=\sum_{i}\delta\left(t-t_{i}^{(k)}\right),$ (11)
where $\delta$ is the Dirac delta function and $t_{i}^{(k)}$ is the time at which the reaction $r_{k}$ occurs for the $i$th time. In this way, the term $\vec{\Lambda}_{J}$ in Equations (5) and (6) manifests as discrete jumps in both $\vec{V}_{F}$ and $\vec{V}_{J}$ at each jump. For the sake of the simulation, the computation of the jump times $t_{i}^{(k)}$ for each instance $i$ of each reaction $r_{k}\in\mathcal{R}_{J}$ is all that is required. The stoichiometric matrix describes the change to each compartment when that jump event occurs.
The rate of reaction $r_{k}$, $\lambda_{k}(t,\vec{V})$, may be state dependent so can change due to the continuous change caused by flows. Figures 2A and 2B depict an example of flowing and jumping compartment (respectively) experiencing a jump event. In this illustration, we note that if $\lambda_{k}$ depends on $v_{2}$ the event time $t_{i}^{k}$ is not sampled from an exponential distribution which requires a constant propensity/rate between events. Formally, the jump times form an inhomogeneous Poisson process. There are multiple ways to sample the jump times $t_{i}^{(k)}$, see [25] for a detailed discussion. We use a variant of the Next Reaction Method [15] (which is an optimised variant of the Doob-Gillespie method [16]) to sample jump times. We first note that the propensity for a jump is dependent only on the instantaneous state $\vec{V}$, and therefore at time $t_{0}$ if there has been $i-1$ jumps associated with reaction ${r}_{k}$, it has no bearing on the distribution of the time $t_{i}^{(k)}$. Therefore, we shall simply denote $t_{i}^{(k)}=t_{k}$ as the next jump time for reaction ${r}_{k}$. The cumulative probability function from which $t_{k}$ is sampled depends on the current time, $t_{0}$, and the evolution of the state variables in time, $\vec{V}(t)$. The cumulative distribution for the time of the next event associated with $r_{k}$ is:
[…]
$J_{k}(t_{k})=-\log(u_{k})-\int_{t_{0}}^{t_{k}}\lambda_{k}(s,\vec{V}(s))\ \mathrm{d}s,$ (13)
noting that $u_{k}$ and $1-u_{k}$ have the same distribution. In general, we cannot solve directly for $t_{k}$, so we solve for it numerically by tracking the value of $J_{k}(t)$ as $\vec{V}$ evolves through flows, jumps and switches.
[…]
To update the jump clock, we require numerical integration of $\lambda_{k}(t,\vec{V}(t))$ forward in time. Fortunately, we also have piece-wise polynomial approximations for $\vec{V}_{F}(t)$ as a result of our numerical treatment of the continuous flows (see Subsection Flow events ‘Flow events’), combined with piece-wise constant values for $\vec{V}_{J}(t)$ which only change when jumps occur. We discuss the numerical integration of the jump clock below.
Jump clock updates
For a given jump reaction ${r}_{k}$, a jump clock is initialised at time $t_{0}$ with $u_{k}\sim\text{Unif}(0,1)$ giving $J_{k}=\log(u_{k}^{-1})$. From time $t$ to $t+\Delta t$, with $t>t_{0}$, Equation (13) tells us the clock ticks down from $J_{k}$ to $J_{k}-\Delta J_{k}$, where $\Delta J_{k}$ is
$\Delta J_{k}=\int_{0}^{\Delta t}\lambda_{k}(t+s,\vec{V}(t+s))\ \mathrm{d}s.$ (14)
In general this integral is intractable, but we can approximate it as follows: consider the Taylor expansion of $\lambda_{k}(t+s,\vec{V}(t+s))$ about $s=0$, which gives
$\Delta J_{k}$ $=\int_{0}^{\Delta t}\lambda_{k}(t,\vec{V})+\dfrac{\partial\lambda_{k}}{\partial\vec{V}}\dfrac{\mathrm{d}\vec{V}^{\intercal}}{\mathrm{d}t}\,s+O(s^{2})\ \mathrm{d}s=\int_{0}^{\Delta t}\alpha+\beta s+O(s^{2})\ \mathrm{d}s,$ (15)
where we write $\alpha=\lambda_{k}(t,\vec{V})$, and $\beta=\dfrac{\partial\lambda_{k}}{\partial\vec{V}}\dfrac{\mathrm{d}\vec{V}^{\intercal}}{\mathrm{d}t}$. Since (i) our numerical evaluation of $\vec{V}$ is piece-wise linear across the time interval, by virtue of the fact that we use a Forward Euler approximation to solving Equation (8), and (ii) $\Delta t$ is small, $\Delta J_{k}$ is approximated to the precision of our algorithm by taking
$\Delta J_{k}$ $\approx\frac{\Delta t}{2}\left(2\alpha+\beta\Delta t\right),$ (16)
noting that since $\vec{V}$ is piece-wise linear across the time interval, the convergence of the integration of the Jump-clock also behaves as $\Delta t$. Importantly, we know that between jumps $\dfrac{\mathrm{d}\vec{V}_{F}}{\mathrm{d}t}$ is given by Equation (8) whilst $\dfrac{\mathrm{d}\vec{V}_{J}}{\mathrm{d}t}=0$. Thus, $\beta=\dfrac{\partial\lambda_{k}}{\partial\vec{V}_{F}}\left(\eta_{FF}\vec{\lambda}_{F}(t,\vec{V})\right)^{\intercal}$ (evaluated at $t$). To calculate the updated jump clock, we compute $J_{k}(t+\Delta t)=J_{k}(t)-\Delta J_{k}$ as the provisional value. We then have two distinct cases: (i) $J_{k}(t)-\Delta J_{k}>0$, then no jump occurred during the interval $(t,t+\Delta t)$ and we have the jump clock $J_{k}(t+\Delta t):=J_{k}(t)-\Delta J_{k}$: (ii) $J_{k}(t)-\Delta J_{k}<0$, then a jump occurred (i.e. a ${r}_{k}$ reaction) during the interval $(t,t+\Delta t)$, that we need to account for in the updated jump clock. In the case (ii) where a jump occurs within the interval $(t,t+\Delta t)$, let $t+\Delta\tau$, where $0<\Delta\tau<\Delta t$, denote the time at which this jump occurs. We can find $\Delta\tau$ by interpolation and solving the following for $\Delta\tau$:
$2\Delta J_{k}-\Delta\tau(2\alpha+\beta\Delta\tau)=0,$ (17)
where $\Delta J_{k}$ is the residual of the jump clock from $t$ to $t+\Delta\tau$. $\Delta\tau$ is then given by
$\Delta\tau=\begin{cases}\frac{\sqrt{\alpha^{2}+2\beta\Delta J_{k}}-\alpha}{\beta},\quad&\beta\neq 0,\\ \frac{\alpha}{\Delta J_{k}},\quad&\beta=0.\end{cases}$ (18)
If there is a jump, rather than using $\Delta t$ to forward compute the flowing compartments in Equation (9) we instead use $\Delta\tau$ to take part of a flow step and then implement the jump after the flow to get the state at time $t+\Delta\tau$. After this, we reinitialise the jump clock $J_{k}$ as described above. This procedure is illustrated in Figure 2C.
Switching events
The way in which we handle switching is one of the main contributions of this manuscript. Switching events occur when variables transition between $\vec{V}_{J}$ and $\vec{V}_{F}$, i.e. between discrete or continuous values. As a consequence, the contents of $\mathcal{R}_{J}$ and $\mathcal{R}_{F}$ may change. There are three types of switching events. The first involves a compartment moving from $\vec{V}_{J}$ to $\vec{V}_{F}$. This transition is straightforward as a new equation is added to Equation (8), and the state $\vec{V}_{F}$ is initialised at the switching time, with the new flowing compartment at $\Omega$. Figure 2D depicts an example of a type 1 switching event. The second type involves a flowing compartment moving from $\vec{V}_{F}$ to $\vec{V}_{J}$. Here, the flowing compartment reaches the switching threshold, $\Omega$, and becomes discrete and no longer follows the flowing dynamics (of Equation (8)). The time at which this occurs is found by continuation, and computation is resumed. Figure 2E depicts an example of a type 2 switching event.
The third type of switching event involves a compartment jumping from $\vec{V}_{F}$ to $\vec{V}_{J}$. In this case, a flowing compartment jumps down to the switching threshold $\Omega$, and becomes discrete. To highlight how the third type of switch event is necessary, we propose considering a simple model where two species, $X$ and $Y$, interact to produce a third species, $Z$, with $X$ also undergoing a death dominated birth-death process. Suppose that the switching threshold is set as $\Omega=1000$, and that $X(0)=1001$, $Y(0)=10$ and $Z(0)=0$. Due to the death reaction of $X$, suppose that within the first (Flow) step, $X$ experiences some decay $\varepsilon<1$. Therefore, $X(t_{1})=1001-\varepsilon>\Omega=1000$ is in a flowing (continuous) regime. Now let’s suppose that in the next instance, $X$ and $Y$ combine to produce one $Z$. At this instance, $X$ must experience some integer loss (which we can assume to be 1, but in general may be any positive value). In this case, producing $Z$ will result in $X(t_{2})<\Omega$ but non integer. Therefore, to account for type three switch events, let $v_{i}$ be the compartment switching from $\vec{V}_{F}$ to $\vec{V}_{J}$ due to a jump. In general, these types of jumps result in $v_{i}$ being non-integer, i.e. $v_{i}\notin\{0,1,\ldots,\Omega\}\cup(\Omega,\infty)$. To ensure the values of $v_{i}$ stay in $\{0,1,\ldots,\Omega\}\cup(\Omega,\infty)$, we add another constraint to the process, initially proposed by [40] [40]. The idea is to randomly round the compartment value up or down to an integer in a way that conserves the average behaviour of the process. Let $\hat{v}_{i}$ be the value of the flowing compartment $v_{i}$ after jumping down across the threshold $\Omega$ but before being rounded into $\{0,1,\ldots,\Omega\}\cup(\Omega,\infty)$. We apply the following rule to round $v_{i}$ after the switch. We take $v_{i}=\lceil\hat{v}_{i}\rceil$ with probability $\hat{v}_{i}-\lfloor\hat{v}_{i}\rfloor$, and otherwise we round down by setting $v_{i}=\lfloor\hat{v}_{i}\rfloor$. This ensures the expected value of the variable after the switch is $\hat{v}_{i}$ as described under the flowing paradigm from which this compartment has come and that the variable remains in the domain $\{0,1,\ldots,\Omega\}\cup(\Omega,\infty)$. Figure 2F depicts an example of a type 3 switching event.
@@ Results
The SIRS model with demography is an extension of the classic susceptible-infectious-removed (SIR) model. In the SIR model susceptible individuals may be infected by contact with infectious individuals; infected individuals eventually cease to be infectious and transition to the removed compartment. The SIRS model with demography extends the SIR model by allowing removed individuals to transition back to being susceptible to infection, for individuals to give birth to new (susceptible) individuals, and to allow for death of individuals (at equal rates). The SIRS model with demography can be written as the following ODE system, with susceptible (S), infectious (I), and recovered (R) individuals:
$\begin{split}\frac{dS}{dt}&=-\beta\frac{IS}{N}+\omega R+\kappa N-\mu S,\\ \frac{dI}{dt}&=\beta\frac{IS}{N}-\gamma I-\mu I,\\ \frac{dR}{dt}&=\gamma I-\omega R-\mu R,\end{split}$ (19)
where $N(t)=S(t)+I(t)+R(t)$ is the total population size, $\beta$ the infection rate, $\gamma$ the recovery rate, $\omega$ the immunity waning rate, $\kappa$ the birth rate, and $\mu$ the death rate. Figure 3A shows a compartmental diagram of the SIRS model with demography.
Figure 3B shows our simulated (via the Doob-Gillespie algorithm) time series of daily noisy measurements of the prevalence of infection. The true parameters used in the simulation are shown in Figure 3C; these values are broadly consistent with existing estimates for influenza or SARS-CoV-2. We assume that these measurements are drawn from a negative binomial distribution where the expected value is equal to the true prevalence of infection and there is a constant (known) dispersion parameter, $k=100$.
@@
Combining the SIRS model with a particle filter [35, 24] we used the first 100 days of the time series to forecast the remaining 150 days. Figure 3B shows the estimated true prevalence across the first 100 days and forecasts the prevalence for the subsequent 150 days. Included in the forecast is a daily estimate of the probability that the pathogen has been eliminated (in the simulation the pathogen was eliminated on day 250). It is important to note that there is an endemic equilibrium in this model, so it is reasonable for the probability to plateau as it does.
[…]
@@
[…]
@@
To further investigate the computational properties of JSF and to compare it with exact (Doob-Gillespie) simulation and (approximate) tau-leaping simulation we used repeated simulations from the SIRS model with demography. We partitioned the simulations into three classes: early stochastic extinction, fade-out after a single epidemic, and sustained transmission (SI Section 2.) As seen in Table SI 3, JSF and exact simulation generate very similar proportions of samples in each of these classes. Two aspects of this system of practical importance are the timing and magnitude of peak infections, and the total number of infections. As can be seen in Figure SI 7 the distribution of peak timing and total number of infections are in good agreement between JSF and exact simulation. For the magnitude of peak infections, there is substantially less variability in the JSF samples than the exact samples, but the proportional difference is small.
With regards to computational efficiency, Figure SI 10B shows the average time required per simulation using JSF, exact simulation (Doob-Gillespie), tau-leaping, and the efficient Tau-hybrid method [33] for the SIRS model for populations of different sizes. For populations of size above about $10^{5}$ (i.e. approximately the size of a small city), JSF is substantially computationally faster than either exact simulation or tau-leaping, there being a difference of at least one order of magnitude in the amount of time needed, and (if a suitable Switching threshold is specified) a constant scaling, irrespective of compartment occupancy, may be obtained. To simplify the interpretation of the results in Figure SI 10B, each of the methods were implemented in the same interpreted language so that the observed differences can be attributed to the algorithm rather than the platform on which it is implemented. We have also included a comprehensive simulation case study of the simple birth-death process (see Section SI 2 of the supplementary information), which exhibits JSf computational efficiency for both birth dominated and death dominated regimes, with comparable results to the exact stochastic process. We also demonstrate JSF’s accuracy for particularly difficult to solve stochastic simulations which exhibit multiple time scales (see Section SI 3 of the supplementary information), where we see that JSF produces summary statistics and trajectories comparable to the exact stochastic process, out performing current hybrid methods.
@@
Figure 4: The TEIRV model describes the within-host dynamics of host cells and virus during infection. The model accounts for delays in virus production post cellular infection and the capability of target cells to enter a refractory state as a defence against infection. A A compartmental diagram of the TEIRV model showing the interactions between: Target cells ($T$), Refractory cells ($R$), cells in the Eclipsed phase of infection ($E$) and Infectious cells ($I$); and the (Total) Virions ($V$). Total Virion particles are utilised as to not distinguish between intra- and extra-cellular Virions. Solid arrows represent the exchange of mass (cells or virions) through compartments, while dashed lines represent the influence of a compartment on a reaction’s rate of occurrence. B The representation of this model as a reaction network with a stoichiometric matrix. This formalises the dependency on different variables indicated by the dashed arrows in the compartmental diagram. C An example trajectory showing the exponential growth and decline in the amount of free virus across the duration of the infection, this process terminates in the virus, eclispse and infected cell populations reaching zero (shortly before time 121212). Reaching zero which is possible in this hybrid continuous/discrete stochastic model but does not occur in a purely ODE based representation.
Understanding viral clearance is important when deciding how long treatment must be maintained. This case study demonstrates how JSF enables inference of virus clearance, which has been computationally challenging to date [44, 12]. In particular, we use a mathematical model — the TEIRV model described below — to study viral clearance using longitudinal data from 6 individuals infected with SARS-CoV-2 [23].
@@
The TEIRV model is an extension of the classic target-infected-virus (TIV) model [38]. In the TIV model target cells may be infected by the virus; before dying, infected cells produce virus; and the virus can degrade or infect remaining target cells. The TEIRV extends the TIV model through the inclusion of an eclipsed compartment, to model the delay between infection of a target cell and the subsequent production of virus, and a refractory compartment, to model heightened antiviral defences of target cells, e.g. through the effects of interferons [23, 6]. To simplify the use of this model a quasi-steady-state approximation for interferon production is employed. This approximation allows us to avoid the need to explicitly model the amount of interferon present.
Figure 4A shows a compartmental diagram of the TEIRV model. Target cells becoming infected by virions at rate $\beta$, which then enter the eclipsed phase. These eclipsed cells then become infectious at rate $k$. Infectious cells are cleared from the population at rate $\delta$, and produce virions at rate $\pi$. Virions are themselves cleared at rate $c$. We note that $V$ is the total number of virion particles, and therefore this model does not distinguish between intra- and extra-cellular virions. As a result, following the infection of a target cell ($T$), no virion particles are lost, in this model. The infectious cells recruit interferon, which cause the target cells to become refractory (and hence protected against infection) at rate $\Phi$. The refractory cells return to a naive state as target cells at rate $\rho$. These assumptions are represented with the following ODE system:
@@
The stoichiometric matrix corresponding to these assumptions is shown in Figure 4B. Figure 4C, which shows a trajectory sampled from the TEIRV model when represented with JSF; the classic boom-bust dynamics of the viral populations can be seen in the exponential growth and decline of the $V$ compartment. Unlike solutions of the ODE model in (20), we see that by time $t=14$, the populations of eclipsed and infected cells, and the virus have gone extinct, indicating a definitive end to the infection.
The basic reproduction number, $\mathcal{R}_{0}$ is a fundamental quantity of epidemiological models. For the TEIRV model, $\mathcal{R}_{0}$ can be calculated from the next-generation matrix [11]:
@@
The longitudinal data contains cycle number (CN) values from nasal samples. The CN values are inversely proportional to the number of virions present. In our analysis we used an existing (empirical) model to link the cycle numbers to the (logarithmic) viral genome load [23]: $\log_{10}V=11.35-0.25\text{CN}$. We assume the observed values are drawn from a normal distribution with mean $\log_{10}V$ and unit standard deviation, and that the values are truncated at the detection limit of 0.65-0.65- 0.65.